If beryllium is heated in the presence of an excess oxygen gas, how much metal is needed to make 2.354g BeO
1
Expert's answer
2019-07-18T07:52:05-0400

m (Be) = m (BeO) × M (Be) / M (BeO) = 2.354 × 9.01 / 25.01 = 0.848 g